# Reinforced Latent Reasoning for LLM-based Recommendation

### v1

@@ Appendix B Compared Methods
    D3 Bao et al. (2024): This is a SOTA LLM-based generative recommendation method. It follows a similar fine-tuning process to BIGRec but differs in its inference strategy. Specifically, it introduces debiasing techniques during inference to enhance the quality of generated recommendations. Additionally, it includes an ensemble design with traditional models; however, we omit this component in our implementation to ensure a fair comparison.
@@
Our method, LatentR3, is compatible with various LLM-based generative recommendation approaches. Accordingly, we implement it on both BIGRec and D3. Notably, existing LLM reasoning methods for recommendation cannot be selected for comparison because: (1) they are not designed for top-N recommendation tasks, and (2) their training relies on explicit CoT supervision, which is difficult to obtain in our setting. Similarly, we exclude general-domain latent reasoning methods for comparison, as they still depend on CoT data for learning. However, in our ablation study, we will include a variant of our method without RL training, which can be viewed as an adaptation of the general latent reasoning method to our setting.
Appendix C Implementation Details
@@
For traditional recommendation models, we use the Adam optimizer and perform grid searches over learning rates 1e-2, 1e-3, 1e-4 and weight decay values 1e-4, 1e-5, 1e-6, and all models are trained using Binary Cross-Entropy loss with randomly sampled negative items. For LLM-based methods, we use Qwen2.5-1.5B Team (2024) as the backbone LLM. Supervised fine-tuning (SFT) is conducted using the AdamW optimizer, with learning rates selected from {3e-3, 3e-4, 3e-5}, and early stopping is applied with a patience of 1. During the reinforcement learning stage, we search learning rates within {1e-5, 1e-4, 5e-4}. All experiments are run on 2 NVIDIA A100 GPUs.
Appendix D Performance on Popular and Unpopular Items
@@
Appendix E Influence of Reasoning Length
@@
[…]
Appendix F Comparison with the Original GRPO
@@
To implement the variant with the original GRPO, after sampling the latent reasoning, we leverage the LLM to generate the final answer, and then exactly match the answer with the ground-truth to compute 0-1 reward. Lastly we leverage the group average for advantage computation. The experiments are conducted on both the Toys and CDs datasets. The results are summarized in Figure 7. In terms of performance, our method achieves results comparable to the variant using original GRPO—slightly outperforming it on the CDs dataset while performing slightly worse on the Toys dataset. In terms of training efficiency, our method is significantly faster, reducing the training cost to approximately 1/5 of that of the original GRPO-based method. These findings validate the rationale and effectiveness of our proposed modifications. Notably, the reported results for the original GRPO are based on full-model tuning during RL, as this setting is necessary to achieve the reported performance. Tuning only the LatentRATT layer, as done in the default LatentR3—results in significantly worse outcomes for GRPO. However, we observe that limiting GRPO to tuning the LatentRATT layer would only slightly reduce training cost (by approximately 50 minutes), indicating that our method remains substantially more efficient.
Appendix G Full Related Work
@@
The recent widespread success of LLMs has sparked growing interest in their application to recommendation systems, giving rise to three main paradigms: (1) in-context learning (Gao et al., 2023; Sun et al., 2024; Wang et al., 2021), (2) agent-based frameworks (Wang et al., 2023; Zhang et al., 2024a, b), and (3) fine-tuning-based methods (Bao et al., 2025; Zhang et al., 2024c). Among these, the fine-tuning-based solution has received the most attention due to its ability to more effectively align LLMs with recommendation objectives, and our work falls into this category. Although significant efforts have been devoted to this direction, most existing methods overlook the role of reasoning in recommendation—particularly latent reasoning, which remains unexplored. Among the limited studies that investigate reasoning, existing works primarily focus on enhancing explicit chain-of-thought (CoT) reasoning through fine-tuning. RecSAVER (Tsai et al., 2024) and ReasoningRec (Bismay et al., 2024) leverage larger LLMs to generate CoT reasoning data for training smaller LLMs, aiming to improve their reasoning capabilities. Reasoning4Rec (Fang et al., 2025) decomposes the reasoning process into multiple steps, using user reviews as proxy supervision. Exp3rt (Kim et al., ) focuses on distilling the reasoning behavior of larger LLMs into smaller models. However, these approaches rely heavily on CoT data, which is costly to obtain and often difficult to ensure in quality. Moreover, the use of explicit CoT reasoning typically incurs higher inference latency. In contrast, our work introduces latent reasoning for LLM-based recommendation, which eliminates the need for CoT supervision and enables more efficient inference. As for latent reasoning, only two concurrent works (Tang et al., 2025; Zhang et al., 2025) have explored it, but neither focuses on LLM-based recommendation, and our approach to latent reasoning is fundamentally different and specifically tailored for LLM-based recommendation.

### v2

@@ Appendix B Compared Methods and Implementation Details
    D3 (Bao et al., 2024): This is a SOTA LLM-based generative recommendation method. It follows a similar fine-tuning process to BIGRec but differs in its inference strategy. Specifically, it introduces debiasing techniques during inference to enhance the quality of generated recommendations. Additionally, it includes an ensemble design with traditional models; however, we omit this component in our implementation to ensure a fair comparison.
@@
Our method, LatentR3, is compatible with various LLM-based generative recommendation approaches. Accordingly, we implement it on both BIGRec and D3. Notably, existing LLM reasoning methods for recommendation cannot be selected for comparison because: (1) they are not designed for top-N recommendation tasks, and (2) their training relies on explicit CoT supervision, which is difficult to obtain in our setting. To align with our setting, we also attempted to learn explicit CoT reasoning directly via RL; however, the results were very poor, as shown in Appendix LABEL:. Similarly, we exclude general-domain latent reasoning methods for comparison, as they still depend on CoT data for learning. However, in our ablation study, we will include a variant of our method without RL training, which can be viewed as an adaptation of the general latent reasoning method to our setting.
@@
$\bullet$ Implementation Details For traditional recommendation models, we use the Adam optimizer and perform grid searches over learning rates 1e-2, 1e-3, 1e-4 and weight decay values 1e-4, 1e-5, 1e-6, and all models are trained using Binary Cross-Entropy loss with randomly sampled negative items. For LLM-based methods, we use Qwen2.5-1.5B Team (2024) as the backbone LLM. Supervised fine-tuning (SFT) is conducted using the AdamW optimizer, with learning rates selected from {3e-3, 3e-4, 3e-5}, and early stopping is applied with a patience of 1. During the reinforcement learning stage, we search learning rates within {1e-5, 1e-4, 5e-4}. All experiments are run on 2 NVIDIA A100 GPUs.
Notably, BIGRec originally uses a grounding-based (matching) method for decoding items. In our implementation, we adopt a constrained decoding approach, as we found it yields better performance. This is supported by the empirical results on the CDs dataset in Table 4.
Table 4: Performance of BIGRec using different decoding methods—(1) Grounding and (2) Constrained—on the CDs dataset.
Decoding method H@5 H@10 N@5 N@10
Grounding Decoding 0.0609 0.0670 0.0573 0.0592
Constrained Decoding 0.0757 0.0929 0.0616 0.0672
Appendix C Learning Explicit Reasoning without COT Supervision via RL
We also experimented with directly optimizing CoT generation via RL, similar to DeepSeek-R1-zero, denoted as “COT-RL.” We compare it with the following methods: Base, COT, and our LatentR3. The results are summarized in Table 5. As shown, both COT and COT-RL outperform the base LLM. However, applying RL to learn CoT reasoning without direct supervision fails to improve performance and can even degrade it. This is likely due to the inherent difficulty of learning long CoT chains: correct sampling is challenging, and the learning process requires considerations beyond accuracy—such as output formatting—which can reduce learning efficiency and lead to reward increases without corresponding performance gains.
Table 5: Performance comparison of explicit reasoning learned via RL (without CoT supervision, denoted by “COT-RL”, ) versus other methods.
Metrics H@5 H@10 N@5 N@10
Base 0.0195 0.0252 0.0148 0.0167
COT 0.0302 0.0406 0.0213 0.0246
COT-RL 0.0297 0.0390 0.0202 0.0232
Ours 0.0934 0.1160 0.0754 0.0826
Appendix D The Analysis of Statistical Significance
In the main result (Table  1), we disabled the sources of randomness to ensure reproducibility. To address your concern, we enabled random sampling during generation and conducted a t-test between our method and the corresponding baseline (BIGRec). When reporting the averaged results, our method remains significantly better, with a p-value < 0.001, as shown in Table 6.
Table 6: T-test results for BIGRec and LatentR3 (applied to BIGRec) on the CDs dataset over five experimental runs. “*” indicates statistical significance with $p<0.01$.
Method H@5 H@10 N@5 N@10
BIGRec 0.0725 0.0902 0.0595 0.0653
+LatentR3 0.0929 0.1131 0.0748 0.0813
p-value 0.00014 0.00067 4.8e-6 1.7e-5
Appendix E Performance on Larger LLM
We have further explored the effectiveness of our method on a larger LLM, Qwen2.5-3B. Considering the high cost of LLM tuning, we focused on comparing our method with BIGRec on only one dataset — CDs. Meanwhile, given the much higher cost of RL tuning for larger LLMs, we apply LoRA-based tuning to control training expenses under the use of larger LLMs. The results are summarized in Table 7. As shown, our method still delivers a more substantial relative improvement on the larger model, achieving an average performance gain of 40.7%.
Table 7: Performance comparison of BIGRec and LatentR3 on the CDs dataset using larger LLMs with LoRA tuning.
Method H@5 H@10 N@5 N@10 Impr.
BIGRec 0.0503 0.0658 0.0389 0.0439 -
+LatentR3 0.0695 0.0903 0.0566 0.0634 40.7%
Appendix F Inference Cost Comparison
Compared to non-reasoning LLM baselines, our method introduces the extra cost of generating latent reasoning tokens. However, since only a few tokens are used (one by default), the added cost is minimal. In contrast, explicit reasoning methods often require a large and uncontrollable number of tokens, leading to significant overhead. We randomly selected 100 samples from each of the four datasets. Inference was performed using a single A100 GPU, with batch size set to 4 and beam size set to 10 (the CoT method set to 1). Each measurement was taken three times and we report the average values, as shown in the Figure 5. Because only a single token is added, and the length of item titles inherently varies, our inference time is almost identical to that of non-reasoning methods. However, explicit CoT methods are far more expensive (COT).
Figure 5: Inference time comparison across non-reasoning (BIGRec), LatentR3 (Ours), and explicit CoT methods (COT).
Appendix G Performance on Popular and Unpopular Items
@@
Figure 6: Performance improvement of LatentR3 over BIGRec on both popular and unpopular items on four dataset.
Appendix H Influence of Reasoning Length
@@
We investigate how length of latent reasoning affects model perfomance by varying the number of generated latent tokens ($K$), with $K\in\text{\{0, 1, 2\}}$. When implementing, a model dedicated to single reasoning token generation was fully trained. We directly perform reinforcement learning to train the 2 reasoning tokens model based on this. Noise sampling was selectively applied only to the last reasoning token, given that the first token’s generative capabilities were already robustly established through the initial training. Experiments are conducted on the Toys and CDs datasets, and the results in terms of NDCG@10 (denoted as N@10), HR@10 (H@10), NDCG@5 (N@5) and HR@5 (H@5) are shown in Figure 7. As shown, increasing the reasoning length improves performance, demonstrating the potential to achieve even better results by increasing the reasoning length. However, the gain from $K\text{=}1$ to $K\text{=}2$ is much smaller than from $K\text{=}0$ to $K\text{=}1$, suggesting that a few latent tokens may be sufficient to capture most of the reasoning information.
Appendix I Comparison with the Original GRPO
@@
To implement the variant with the original GRPO, after sampling the latent reasoning, we leverage the LLM to generate the final answer, and then exactly match the answer with the ground-truth to compute 0-1 reward. Lastly we leverage the group average for advantage computation. The experiments are conducted on both the Toys and CDs datasets. The results are summarized in Figure 8. In terms of performance, our method achieves results comparable to the variant using original GRPO—slightly outperforming it on the CDs dataset while performing slightly worse on the Toys dataset. In terms of training efficiency, our method is significantly faster, reducing the training cost to approximately 1/4 of that of the original GRPO-based method. These findings validate the rationale and effectiveness of our proposed modifications. Notably, the reported results for the original GRPO are based on full-model tuning during RL, as this setting is necessary to achieve the reported performance. Tuning only the LatentRATT layer, as done in the default LatentR3—results in significantly worse outcomes for GRPO. However, we observe that limiting GRPO to tuning the LatentRATT layer would only slightly reduce training cost (by approximately 50 minutes), indicating that our method remains substantially more efficient.
Appendix J Full Related Work
@@
The recent widespread success of LLMs has sparked growing interest in their application to recommendation systems, giving rise to three main paradigms: (1) in-context learning (Gao et al., 2023; Sun et al., 2024; Wang et al., 2021), (2) agent-based frameworks (Wang et al., 2023; Zhang et al., 2024a; b), and (3) fine-tuning-based methods (Bao et al., 2025; Zhang et al., 2024c). Among these, the fine-tuning-based solution has received the most attention due to its ability to more effectively align LLMs with recommendation objectives, and our work falls into this category. Although significant efforts have been devoted to this direction, most existing methods overlook the role of reasoning in recommendation—particularly latent reasoning, which remains unexplored. Among the limited studies that investigate reasoning, existing works primarily focus on enhancing explicit chain-of-thought (CoT) reasoning through fine-tuning. RecSAVER (Tsai et al., 2024) and ReasoningRec (Bismay et al., 2025) leverage larger LLMs to generate CoT reasoning data for training smaller LLMs, aiming to improve their reasoning capabilities. Reasoning4Rec (Fang et al., 2025) decomposes the reasoning process into multiple steps, using user reviews as proxy supervision. Exp3rt (Kim et al., ) focuses on distilling the reasoning behavior of larger LLMs into smaller models. However, these approaches rely heavily on CoT data, which is costly to obtain and often difficult to ensure in quality. Moreover, the use of explicit CoT reasoning typically incurs higher inference latency. In contrast, our work introduces latent reasoning for LLM-based recommendation, which eliminates the need for CoT supervision and enables more efficient inference. As for latent reasoning, only two concurrent works (Tang et al., 2025; Zhang et al., 2025) have explored it, but neither focuses on LLM-based recommendation, and our approach to latent reasoning is fundamentally different and specifically tailored for LLM-based recommendation.
@@
Appendix K The Use of Large Language Models
The use of Large Language Models (LLMs) is an integral part of our methodology. We employ LLMs primarily to assist in the refinement of textual content, including the optimization of language and clarity in the writing of this paper. A detailed description of how the LLM was utilized within our methodological framework is provided in Section 3. It is important to clarify that while LLMs were used as tools to support drafting and editing, all research ideas, conceptual design, data analysis, and final decision-making were conducted entirely by the human authors.